# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v1

@@ III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
This problem can be solved using non-linear solvers or the successive convex approximation optimization method. Both of these are computationally expensive for large network deployment. Therefore, to solve this problem with low computation cost, we employ gradient based proximal gradient ascent (PGA) method [22, 23]. We first reformulate the problem to be unconstrained except for the box constraint (19). This is achieved by moving the other constraints into the objective function via a penalty method. The reformulated problem is:
$\underset{\bm{\delta}_{m}}{\max~}$ $f(\bm{\delta}_{m}){=}\sum\limits_{t=1}^{T}\log_{2}(1+\text{SINR}^{\text{G-PFZF}}_{mt})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\!\Bigg),$ (22a)
@@
$\bm{{\delta}}^{(j+1)}_{m}=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)=\min\!\left(1,\max\!\left(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)\right).$ (23)
The components of gradient $\nabla f(\bm{\delta}_{m})=\Big[\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{m1}},\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{m2}},\dots,\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mL_{p}}}\Big]^{\intercal}$ can be written as
@@
$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{u}_{k}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{m}-\sum\limits_{k\in\mathcal{P}_{i_{t}}}p^{u}_{k}\gamma_{mk},&\text{if }\ i_{t}=i,\\ -\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\sum\limits_{k\in\mathcal{P}_{i}}p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},&\text{if }\ i_{t}\neq i,\end{cases}$ (26)
@@
5:    Compute $\bm{{\delta}}^{(j+1)}_{m}$ using (23)
@@
Convergence and Computation Cost Analysis: Since the feasible set defined by the box constraint (19) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proof is given in Appendix -B) with constant $L>0$, one should choose $\alpha\in(0,1/L]$ to ensure that the gradient step remains within the stability region of the PGA method [22, 23]. In numerical simulations, small step size parameter $\alpha$ has been observed to achieve convergence for Algorithm 1.
@@
Note 2: To reduce the computational cost, we use pilot-based grouping instead of UE-based grouping, as discussed in [19, 13]. In UE-based grouping, each AP would have $T$ binary variables (one per UE), leading to $TM$ total variables across the network. In the proposed pilot-based grouping, each AP has $L_{p}$ binary variables, leading to $ML_{p}$ total variables. Since $L_{p}\ll T$ in dense deployments, this significantly reduces the computational cost.
@@ III-B Generalized Protected Weak Partial Full-Pilot Zero Forcing Combining Scheme
In the G-PFZF scheme, the combiner for the UEs assigned to a strong pilot is designed only to suppress interference only from other UEs that are also on strong pilots. This intra-group suppression leaves UEs on weak pilots vulnerable to dominant interference from the strong UEs.
The proposed G-PWPFZF scheme introduces a more comprehensive, protective strategy. In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that use a strong pilot. This is achieved by projecting the combining vectors for weak UEs onto the orthogonal complement of the subspace spanned by the strong UEs’ effective channels. Consequently, the G-PWPFZF scheme provides universal protection against the most significant sources of interference, dramatically improving performance for far users on weak pilots, at cost of higher computation cost of designing the vector for weak pilot UEs.
In the G-PWPFZF scheme, the strong pilots are assigned with the vectors in the same way as in (7). The combining vector allocated to the pilot $i_{t}\in\mathcal{W}_{m}$ corresponding to the AP $m$ is simply the normalized MR combining vector, as in (8), with a projection matrix and is given by:
${\mathbf{v}}^{\text{PMR}}_{mi_{t}}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi_{t}}}}\mathbf{B}_{m}\mathbf{\bar{G}}_{m}e_{i_{t}},$ (27)
where $\mathbf{B}_{m}$ is a projection matrix that projects the received signal onto the orthogonal complement of the subspace spanned by the effective channels of all strong pilots. It is defined as [19]:
@@
The expected value of the effective channel gain for the UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}_{mk}}\}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (28)
Furthermore, the second moment of the effective channel gain is given by:
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (29)
@@
By using the combining vector (7) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by:
@@
The derivation of SINR expression is given in Appendix -C.
Local Weights Design for G-PWPFZF Scheme: The design of local weights for the G-PWPFZF scheme follows a similar design to that of the G-PFZF scheme, aiming to maximize the local contribution to the final SINR while suppressing interference in a fully distributed manner. Here too the local weights are same as the local SINR.
@@
Adaptive Pilot Grouping for G-PWPFZF Combining: This subsection discusses the local optimization framework for pilot grouping in the G-PWPFZF scheme. The objective and constraint structure remain identical to those of the G-PFZF scheme. The sole distinction lies in the expression for the local SINR, which forms the foundation of the utility function $f(\bm{\delta}_{m})$ for pilot-grouping. The local SINR for the UE $t$ at the AP $m$ for G-PWPFZF scheme, given in (31) after substituting $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$, can be rewritten as:
[…]
Consequently, the optimization algorithm (Algorithm 1) and the overall structure of the gradient $\nabla f(\bm{\delta}_{m})$ remain unchanged. The impact of the different SINR is confined solely to the calculation of the partial derivatives $\frac{\partial S_{mt}}{\partial\delta_{mi}}$ and $\frac{\partial I_{mt}}{\partial\delta_{mi}}$ within the gradient. For the G-PWPFZF SINR in (32), these derivatives are:
@@ IV Numerical Simulations
This section demonstrates the influence of the proposed improvements on the SE performance across multiple distributed combining strategies and local weights. We consider a network with $M$ APs and $T$ UEs, uniformly distributed in a square area of size $1\times 1$ km2. The parameters are mentioned in Table I. Large-scale fading coefficients (LSFCs) are generated according to the model in [3], incorporating shadow fading. Pilots are assigned using the scalable pilot assignment strategy from [3], ensuring efficient reuse under limited pilot resources. All APs are assumed to operate at full transmit power during both pilot training and data transmission. For the baseline PFZF and PWPFZF schemes, UE grouping is performed using a $90\%$ of total LSFCs received at the AP, which we identified using numerical simulation as providing the best performance. To ensure statistical reliability, all reported results are averaged over 500 independent simulation realizations.
@@ IV-A Computation cost and Fronthaul cost
In this subsection, we analyze the computational cost and fronthaul signaling cost. The costs are computed following the approach outlined in [16, 19, 15, 10], where only multiplication and division operations are counted per coherence block, as these dominate the computational load.
@@
Decoding Type Fronthaul Cost (Complex Scalars) Combining Weights Computation Cost
o-LSFD [16, 15] $L_{u}M+\frac{3MT+M}{2}$ $\left(\frac{M^{2}+M}{2}\right)T+\frac{M^{3}-M}{3}+M^{2}$
Proposed Decentralized Decoding $L_{u}M$ $\left(\frac{M^{2}+M}{2}\right)T+M+M^{2}$
Table II compares the fronthaul and computational costs for optimal o-LSFD and the proposed decentralized decoding architecture. In conventional o-LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead of $L_{u}M+\frac{3MT+M}{2}$ complex scalars per UE per coherence block. Furthermore, o-LSFD requires computationally expensive matrix operations such as matrix inversion at the CPU that scale as $\mathcal{O}(TM^{2}+M^{3})$ per UE, making it impractical for large-scale deployments.
The proposed decentralized decoding architecture eliminates these bottlenecks by avoiding centralized LSFD entirely. Each AP independently computes and applies interference-suppressing local weights, reducing the fronthaul load to only $L_{u}M$ complex scalars. The computational cost is similarly streamlined to $\mathcal{O}(TM^{2})$ by eliminating the cubic $M^{3}$ term associated with large matrix inversions. This combination of reduced fronthaul overhead and scalable computation enables practical deployment in dense network scenarios.
Figure 2: Computational cost of weights calculation versus number of UEs, showing the significant computation cost reduction of the proposed decentralized decoding approach compared to o-LSFD.
Figure 3: Fronthaul cost versus number of UEs, demonstrating the reduction achieved by the proposed decentralized decoding architecture by eliminating channel estimate transmission.
As demonstrated in Fig. 2, the proposed decentralized decoding weights computation provides substantial computation cost reduction compared to o-LSFD, particularly as the number of UEs increases. Similarly, Fig. 3 shows that eliminating channel estimate transmission reduces fronthaul overhead.
@@
Figure 4: Computational cost of combining vectors versus number of UEs, illustrating the scalability advantage of generalized schemes (G-PFZF, G-PWPFZF) over traditional approaches.
The computational cost per AP per coherence block for calculating the combining vectors is detailed in Table III. The primary distinction between the traditional and proposed generalized schemes lies in their scaling behavior relative to the number of users. The traditional PFZF and PWPFZF schemes exhibit a computational cost that scales linearly with the total number of user equipments, $T$, as in by the $AT$ term for PFZF and PWPFZF. This dependence creates a significant computational bottleneck in densed networks with massive connectivity. In contrast, the proposed G-PFZF and G-PWPFZF schemes achieve a fundamental scalability advantage by scaling with the number of pilots, $L_{p}$, rather than the number of UEs. This is reflected in the $AL_{p}$ term for G-PFZF and G-PWPFZF, which is a fixed system parameter independent of the user population. Fig. 4 further validates the scalability benefits of the generalized schemes, where computational cost saturates once the number of UEs exceeds the available pilot sequences, unlike traditional schemes that continue to scale linearly with user density.

### v2

@@ III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
This problem can be solved using non-convex solvers or the successive convex approximation optimization method. Both of these are computationally expensive for large network deployment as non-convex solvers. Therefore, to solve this problem with low computation cost, we employ gradient based proximal gradient ascent (PGA) method [22, 23]. We first reformulate the problem to be unconstrained except for the box constraint (19). The reformulated problem is:
$\underset{\bm{\delta}_{m}}{\max~}$ $f(\bm{\delta}_{m}),$ (22a)
@@
where
$\begin{split}f(\bm{\delta}_{m})&=\!\!\sum\limits_{t=1}^{T}\log_{2}(1+\frac{S_{mt}}{I_{mt}})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\Bigg),\end{split}$
@@
$\begin{split}\bm{{\delta}}^{(j+1)}_{m}&=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)=\min\!\left(1,\max\!\left(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)\right).\end{split}$ (23)
The components of gradient $\nabla f(\bm{\delta}_{m})$ can be written as
@@
$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\ \Big(1+\!\!\!\!\!\!\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\!\!\!\!\!\!\!\delta_{mj}\Big)\gamma_{m}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\gamma_{mk},\ \ \text{if }i_{t}=i,\\ \vskip 5.69054pt-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},\ \ \text{if }i_{t}\neq i,\end{cases}$ (26)
@@
5:    Compute:
$\bm{{\delta}}^{(j+1)}_{m}{=}\min\!\big(1,\max\!\big(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\big)\big)$
@@
Convergence and Computation Cost Analysis: Since the feasible set defined by the box constraint (19) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proved in Appendix -B) with constant $L>0$, one should choose $\alpha\in(0,1/L]$ to ensure the convergence of the PGA method [22, 23]. In numerical simulations, small step size parameter $\alpha$ has been observed to achieve convergence for Algorithm 1.
@@
Note 2: To reduce the computational cost, we use pilot-based grouping instead of UE-based grouping as in [19, 13]. In UE-based grouping, each AP would have $T$ binary variables (one per UE), leading to $TM$ total variables across the network. In the proposed pilot-based grouping, each AP has $L_{p}$ binary variables, leading to $ML_{p}$ total variables. Since $L_{p}\ll T$ in dense deployments, this significantly reduces the computational cost.
@@ III-B Generalized Protected Weak Partial Full-Pilot Zero Forcing Combining Scheme
In the G-PFZF scheme, the combiner for the UEs assigned to a strong pilot is designed only to suppress interference from other UEs that are also on strong pilots. This intra-group suppression leaves UEs on weak pilots vulnerable to dominant interference from the strong UEs.
The proposed G-PWPFZF scheme introduces a more comprehensive, protective strategy. In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that have a strong pilot. This is achieved by projecting the combining vectors for weak UEs onto the orthogonal complement of the subspace spanned by the strong UEs’ effective channels. Consequently, the G-PWPFZF scheme provides universal protection against the most significant sources of interference, improving performance for users on weak pilots, at the cost of higher computation in comparison to G-PFZF scheme due to projection matrix computation at the AP.
In the G-PWPFZF scheme, the strong pilot $i{\in}\mathcal{S}_{m}$ are assigned with the combining vectors ${\mathbf{v}}^{\text{LFZ}}_{mi}$ as in (7). The combining vector allocated to pilot $i{\in}\mathcal{W}_{m}$ at AP $m$ is the projected MR (PMR) combining vector, and is given by:
${\mathbf{v}}^{\text{PMR}}_{mi}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi}}}\mathbf{B}_{m}\mathbf{\bar{G}}_{m}e_{i},$ (27)
where $\mathbf{B}_{m}$ is a matrix that projects the received signal onto the orthogonal complement of the subspace spanned by the channels of all strong pilots. It is defined as [19]:
@@
The expected value of the effective channel gain for UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}{\mathbf{\hat{g}}_{mk}}\}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (28)
Furthermore
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\neq i_{t},\end{cases}$ (29)
@@
By using the combining vector (7) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by
@@
The derivation of the SINR expression is given in Appendix -C.
Local Weights Design for G-PWPFZF Scheme: The design of local weights for the G-PWPFZF scheme follows a similar design to that of the G-PFZF scheme, aiming to maximize the local contribution to the final SINR while suppressing interference in a fully decentralized manner. Here too the local weights are the same as the local SINR.
@@
Adaptive Pilot Grouping for G-PWPFZF Combining: This subsection discusses the local optimization framework for pilot grouping in the G-PWPFZF scheme. The objective and constraint structure remain identical to those of the G-PFZF scheme. The sole distinction lies in the expression for the local SINR, which forms the foundation of the utility function $f(\bm{\delta}_{m})$ for pilot-grouping. The local SINR for UE $t$ at AP $m$ for G-PWPFZF scheme is given in (31). After substituting $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$ in (31):
$S_{mt}=p^{u}_{t}\Big(A-\sum\limits_{i=1}^{L_{p}}\delta_{mi}\Big)\gamma_{mt}$
$I_{mt}=\!\!\!\!\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!p^{u}_{k}\Big(\!A-\!\!\sum\limits_{i=1}^{L_{p}}\delta_{mi}\!\Big)\gamma_{mk}+\!\sum\limits_{k=1}^{K}\!p^{u}_{k}(\beta_{mk}{-}\delta_{mi_{k}}\gamma_{mk}){+}1.$
Consequently, the optimization algorithm (Algorithm 1) and the overall structure of the gradient $\nabla f(\bm{\delta}_{m})$ remain unchanged. The impact of the different SINR is confined solely to the calculation of the partial derivatives $\frac{\partial S_{mt}}{\partial\delta_{mi}}$ and $\frac{\partial I_{mt}}{\partial\delta_{mi}}$ within the gradient. For the G-PWPFZF scheme, these derivatives are:
@@ IV Performance Evaluation
This section demonstrates the gains of the our proposed decentralized decoding and combining schemes in terms of the sum SE, per user SE and computation cost. We consider a network with $M$ APs and $T$ UEs, uniformly distributed in a square area of size $1\times 1$ km2. The relevant system parameters are mentioned in Table I. The LSFCs are generated according to the model in [3], incorporating shadow fading. Pilots are assigned using the scalable pilot assignment strategy from [3], ensuring efficient reuse under limited pilot resources. All APs are assumed to operate at full transmit power during both pilot training and data transmission. For the baseline PFZF and PWPFZF schemes, UE grouping is performed using a $90\%$ of total LSFCs received at the AP, are identified using numerical simulation, to provide the best performance. To ensure statistical reliability, all reported results are averaged over 500 independent network realizations.
@@ IV-A Computation cost and Fronthaul cost
In this subsection, we analyze the computational and fronthaul signaling costs of our proposed decentralized decoding and combining schemes. The costs are computed following the approach outlined in [16, 19, 15, 10], where only multiplication and division operations are counted per coherence block, as these dominate the computational load.
@@
Decoding Type Fronthaul Cost Computation Cost
LSFD [16, 15] $L_{u}M+\frac{3MT+M}{2}$ $\left(\frac{M^{2}+M}{2}\right)T+\frac{M^{3}-M}{3}+M^{2}$
Proposed Decoding $L_{u}M$ $\left(\frac{M^{2}+M}{2}\right)T+M+M^{2}$
Table II compares the fronthaul and computational costs for optimal LSFD and the proposed decentralized decoding architecture. In conventional LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead of $L_{u}M+\frac{3MT+M}{2}$ complex scalars per UE per coherence block. Furthermore, LSFD requires computationally expensive matrix operations such as matrix inversion at the CPU that scale as $\mathcal{O}(TM^{2}+M^{3})$ per UE, making it impractical for large-scale network deployments.
The proposed decentralized decoding architecture eliminates these bottlenecks by avoiding centralized LSFD entirely. Each AP independently computes and applies interference-suppressing local weights, reducing the fronthaul load to only $L_{u}M$ complex scalars. The computational cost is similarly streamlined to $\mathcal{O}(TM^{2})$ by eliminating the cubic $M^{3}$ term associated with large matrix inversions required in the LSFD architecture due to centralized AP coordination. This combination of reduced fronthaul overhead and scalable computation enables practical deployment in dense network scenarios.
Figure 2: Comparison of computational and fronthaul costs for different decoding and combining schemes.
As demonstrated in Fig. 2a, the proposed decentralized decoding weights computation provides substantial computation cost reduction compared to LSFD, particularly as the number of UEs increases. Similarly, Fig. 2b shows that eliminating channel estimate transmission reduces fronthaul overhead.
@@
The computational cost per AP per coherence block for calculating the combining vectors is detailed in Table III. The primary distinction between the traditional and proposed generalized schemes lies in their scaling behavior relative to the number of users. The traditional PFZF and PWPFZF schemes exhibit computational costs that scale linearly with the total number of user equipments, $T$, as in the $AT$ term for PFZF and PWPFZF. This dependence creates a significant computational bottleneck in densed networks with massive connectivity. In contrast, the proposed G-PFZF and G-PWPFZF schemes achieve a fundamental scalability advantage by scaling with the number of pilots, $L_{p}$, rather than the number of UEs. This is reflected in the $AL_{p}$ term for G-PFZF and G-PWPFZF, which is a fixed system parameter independent of the user population. Fig. 2c further validates the scalability benefits of the generalized schemes, where computational cost saturates once the number of UEs exceeds the available pilot sequences, unlike traditional schemes that continue to scale linearly with user density.